OpenStudy (anonymous):

Haldifor Steel Products used 400 sheets of steel to make heavy-duty shelving. Each sheet weighed 32.8 kilograms. In making the shelving 2% of the steel used became scrap. How much did the scrap weigh in kilograms?

OpenStudy (daniellelovee):

400x32.8kg=? and then find the 2 percent of ? and substract it from the total ?

OpenStudy (anonymous):

13120 kg

OpenStudy (daniellelovee):

ok now 13120/100=x/2

OpenStudy (anonymous):

so the 2% of 13120 is 262.4

OpenStudy (daniellelovee):

ok if im honest with you this seems more difficult than when I first read it but I would substract 400 minus that value but I think I did this wrong if something you can ask the @mathmate tomorrow

OpenStudy (daniellelovee):

he/she is very good at helping understand difficult problems

OpenStudy (daniellelovee):

sorry I wasnt of any help :/

OpenStudy (anonymous):

Thanks thats ok

OpenStudy (jhannybean):

\[\sf \text{400 sheets of steel} \cdot \frac{\text{32.8 kilograms}}{\text{1 sheet}} = 13120~\text{ kilograms of steel}\]\[\sf 13120~~ \text{kilograms of steel} \cdot \frac{2}{100}=\text{262.4 kilograms of steel}\]

OpenStudy (mathmate):

@keegq I think @daniellelovee 's doubt was what to do next. In fact, she has correctly helped you to get the quantity (weight) of scrap, which @jhannybean has recalculated for you. Since we are told "2% of the steel used became scrap", so the 2% value of 262.4 kg you calculated is the correct answer. If you have doubts how to get that, review jhannybean's work and post if you have questions.